# JEL classification: C31, J24

### v1

@@ 3.3 Simulation Study
$\lambda=0.7$ $0.651$ $0.038$ $0.422$ $0.021$ $0.536$ $0.023$ $0.701$ $0.018$
$\tilde{\beta}_{1}=1$ $2.995$ $0.337$ $0.976$ $0.087$ $0.992$ $0.059$ $1.001$ $0.047$
$\tilde{\beta}_{2}=1.5$ $3.069$ $0.730$ $1.437$ $0.149$ $1.490$ $0.107$ $1.499$ $0.083$
$\tilde{\gamma}_{1}=5$ $4.026$ $0.374$ $3.464$ $0.310$ $4.308$ $0.163$ $5.003$ $0.098$
$\tilde{\gamma}_{2}=-3$ $-9.189$ $0.504$ $-6.719$ $0.344$ $-3.365$ $0.442$ $-2.997$ $0.172$
$\sigma^{2}_{\epsilon}=8$ $9.135$ $8.653$ $9.122$ $5.216$
$\sigma^{2}_{\eta}=15$ $26.975$ $10.576$ $15.329$ $6.024$
$\rho=0.4$ $0.198$ $0.661$ $0.459$ $0.511$
  • Model 1 disregards school heterogeneity (i.e., $c_{s}$ is constant across schools and $\alpha_{s}=0$), while Model 2 accounts for it (with $\alpha_{s}=0$). Model 3 includes a single fixed effect per school with a dummy variable capturing isolated students (i.e., $\alpha_{s}$ may not be zero but is constant across schools, whereas $c_{s}$ varies). Model 4 is our structural model.
@@ 4.1 Data
After removing observations with missing data, the sample used for our empirical analysis encompasses 68,430 students from 141 schools. The number of students per school ranges widely from 18 to 2,027, with an average of 485 per school. On average, each student reports having 3.4 friends (1.6 male friends and 1.9 female friends). Moreover, there are 14,900 (22%) students who have no peers (isolated students), including 7,655 (11%) who are not fully isolated, that is, they are nominated by others. Only 1% of students nominate 10 friends. This suggests that the top coding issue resulting from students’ inability to nominate more than 10 friends (Griffith, 2022) is not a serious concern here.
The dependent variable, GPA, is the average grade across four subjects: mathematics, science, English/language arts, and history or social science. GPA is calculated on the basis of students’ grades in these four subjects, which we recoded as follows: A = 4, B = 3, C = 2, and D = 1. In our analysis, we include controls for several potential factors that may influence GPA (Duncan et al., 2001; Lin, 2010). These factors include sex, age, Hispanic ethnicity, race, living arrangements (whether the student lives with both parents), duration of attendance at the current school, participation in school clubs, mother’s education level, and mother’s profession. We also control for contextual variables associated with the student’s social network by including the average of friends’ control variables.

### v2

@@ 3.3 Simulation Study
$\lambda=0.7$ $0.420$ $0.026$ $0.600$ $0.024$ $0.701$ $0.022$
$\beta_{0,1}=1$ $0.988$ $0.069$ $0.995$ $0.052$ $1.000$ $0.048$
$\beta_{0,2}=1.5$ $1.446$ $0.127$ $0.995$ $0.052$ $1.500$ $0.084$
$\gamma_{0,1}=5$ $4.426$ $0.192$ $4.769$ $0.108$ $4.997$ $0.092$
$\gamma_{0,2}=-3$ 5.423-5.423 $0.394$ $4.769$ $0.108$ 3.005-3.005 $0.163$
10% of Isolated Nodes
DGP A
$\lambda=0.7$ $0.699$ $0.013$ $0.699$ $0.013$ $0.699$ $0.020$
$\beta_{0,1}=1$ $1.002$ $0.045$ $1.001$ $0.045$ $1.001$ $0.046$
$\beta_{0,2}=1.5$ $1.503$ $0.084$ $1.504$ $0.084$ $1.503$ $0.085$
$\gamma_{0,1}=5$ $5.004$ $0.091$ $5.004$ $0.092$ $5.005$ $0.093$
$\gamma_{0,2}=-3$ 3.007-3.007 $0.099$ 3.006-3.006 $0.131$ 3.007-3.007 $0.165$
DGP B
$\lambda=0.7$ $0.465$ $0.032$ $0.699$ $0.013$ $0.699$ $0.020$
$\beta_{0,1}=1$ $0.983$ $0.082$ $1.001$ $0.045$ $1.001$ $0.046$
$\beta_{0,2}=1.5$ $1.425$ $0.154$ $1.504$ $0.084$ $1.503$ $0.085$
$\gamma_{0,1}=5$ $4.220$ $0.234$ $5.004$ $0.092$ $5.005$ $0.093$
$\gamma_{0,2}=-3$ 6.844-6.844 $0.490$ 3.006-3.006 $0.131$ 3.007-3.007 $0.165$
DGP C
$\lambda=0.7$ $0.421$ $0.021$ $0.567$ $0.023$ $0.699$ $0.020$
$\beta_{0,1}=1$ $0.980$ $0.075$ 50.032-50.032 $5.157$ $1.001$ $0.046$
$\beta_{0,2}=1.5$ $1.435$ $0.138$ $0.993$ $0.055$ $1.503$ $0.085$
$\gamma_{0,1}=5$ $4.047$ $0.235$ $1.496$ $0.096$ $5.005$ $0.093$
$\gamma_{0,2}=-3$ 6.230-6.230 $0.339$ $4.602$ $0.129$ 3.007-3.007 $0.165$
22% of Isolated Nodes (Add Health)
DGP A
$\lambda=0.7$ $0.700$ $0.013$ $0.700$ $0.013$ $0.699$ $0.020$
$\beta_{0,1}=1$ $1.000$ $0.046$ $1.000$ $0.046$ $1.000$ $0.046$
$\beta_{0,2}=1.5$ $1.502$ $0.086$ $1.502$ $0.086$ $1.502$ $0.086$
$\gamma_{0,1}=5$ $5.006$ $0.096$ $5.006$ $0.100$ $5.004$ $0.102$
$\gamma_{0,2}=-3$ 3.004-3.004 $0.090$ 3.003-3.003 $0.127$ 2.998-2.998 $0.183$
DGP B
$\lambda=0.7$ $0.480$ $0.034$ $0.700$ $0.013$ $0.699$ $0.020$
$\beta_{0,1}=1$ $0.975$ $0.097$ $1.000$ $0.046$ $1.000$ $0.046$
$\beta_{0,2}=1.5$ $1.419$ $0.172$ $1.502$ $0.086$ $1.502$ $0.086$
$\gamma_{0,1}=5$ $3.699$ $0.313$ $5.006$ $0.100$ $5.004$ $0.102$
$\gamma_{0,2}=-3$ 7.757-7.757 $0.518$ 3.003-3.003 $0.127$ 2.998-2.998 $0.183$
DGP C
$\lambda=0.7$ $0.415$ $0.021$ $0.525$ $0.024$ $0.699$ $0.020$
$\beta_{0,1}=1$ $0.970$ $0.081$ 43.385-43.385 $5.077$ $1.000$ $0.046$
$\beta_{0,2}=1.5$ $1.435$ $0.153$ $0.986$ $0.058$ $1.502$ $0.086$
$\gamma_{0,1}=5$ $3.442$ $0.307$ $1.492$ $0.110$ $5.004$ $0.102$
$\gamma_{0,2}=-3$ 6.718-6.718 $0.330$ $4.291$ $0.166$ 2.998-2.998 $0.183$
  • Model 1 accounts for a single fixed effect per school. Model 2 includes a single fixed effect per school with a dummy variable capturing isolated students. Model 3 is our structural model.
@@ 4.1 Data
After removing observations with missing data, the sample used for our empirical analysis encompasses 68,430 students from 141 schools. The number of students per school ranges widely from 18 to 2,027, with an average of 485 per school. On average, each student reports having 3.4 friends (1.6 male friends and 1.9 female friends). Moreover, there are 14,900 (22%) students who have no peers (isolated students), including 7,655 (11%) who are not fully isolated, that is, they are nominated by others. Only 1% of students nominate 10 friends, suggesting that the top coding issue resulting from students’ inability to nominate more than 10 friends is not a serious concern here.
The dependent variable, GPA, is the average grade across four subjects: mathematics, science, English/language arts, and history or social science. GPA is calculated on the basis of students’ grades in these four subjects, which we recoded as follows: A = 4, B = 3, C = 2, and D = 1. In our analysis, we include controls for several potential factors that may influence GPA. These factors include sex, age, Hispanic ethnicity, race, living arrangements (whether the student lives with both parents), duration of attendance at the current school, participation in school clubs, mother’s education level, and mother’s profession. We also control for contextual variables associated with the student’s social network by including the average of friends’ control variables.